# A generative, multi-output, latent-variable model for astronomical data with noisy labels

### v1

@@ Abstract
The large volume of spectroscopic data available now and from near-future surveys will enable high-dimensional measurements of stellar parameters and properties. Current methods for determining stellar labels from spectra use physics-driven models, which are computationally expensive and have limitations in their accuracy due to simplifications. While machine learning methods provide efficient paths toward emulating physics-based pipelines, they often do not properly account for uncertainties and have complex model structure, both of which can lead to biases and inaccurate label inference. Here we present Lux: a data-driven framework for modeling stellar spectra and labels that addresses prior limitations. Lux is a generative, multi-output, latent variable model framework built on JAX for computational efficiency and flexibility. As a generative model, Lux properly accounts for uncertainties and missing data in the input stellar labels and spectral data and can either be used in probabilistic or discriminative settings. Here, we present several examples of how Lux can successfully emulate methods for precise stellar label determinations for stars ranging in stellar type and signal-to-noise from the APOGEE surveys. We also show how a simple Lux model is successful at performing label transfer between the APOGEE and GALAH surveys. Lux is a powerful new framework for the analysis of large-scale spectroscopic survey data. Its ability to handle uncertainties while maintaining high precision makes it particularly valuable for stellar survey label inference and cross-survey analysis, and the flexible model structure allows for easy extension to other data types.
@@ 1 Introduction
In an attempt to circumvent these requirements, in the last decade there has been a push to use data-driven methods to determine stellar parameters and element abundances of stars using linear regression (e.g., Ness et al., 2015; Casey et al., 2016) or machine-learning methods (e.g., Ting et al., 2019; Ciuca & Ting, 2022; Andrae et al., 2023a; Buck & Schwarz, 2024; Różański et al., 2024), that are more suited to deal with high-volume data. These methods fall under the umbrella of “emulator” or “label transfer” approaches depending on the training and testing data employed, and in essence are used to: 1) train a model on a set of (trustworthy) input stellar spectra and stellar labels; 2) optimize a set of (latent) model parameters; 3) use the trained model to predict stellar labels for some catalog data. While the functional form of the model may vary between these approaches (e.g., quadratic, neural-network, etc…), the process is still the same in practice. These models have proven extremely successful in delivering accurate and precise stellar labels in a fast and cost-effective manner (e.g., Ness et al., 2016; Ho et al., 2017b, a; Xiang et al., 2017, 2019; Andrae et al., 2023b; Li et al., 2024; Guiglion et al., 2024; Ness et al., 2024). However, they also come with some limitations. For example, these models assume the input training stellar labels are ground-truth (i.e., no uncertainties are taken into account), and are not able to train a model with missing label data (for example, stars that have $T_{\mathrm{eff}}$ information but no [Fe/H]). As a result, many good data are not used in the training step, which hinders the stellar label regime that can be probed. This restriction also limits the ability to perform a two-way label transfer between two spectroscopic data sets, as typically stars will have a set of labels from one survey but not the other.
@@
In Section 2 we introduce the data and samples used. In Section 3 we describe the framework of the Lux model, introduce the likelihood function, explain the routine for employing the model, and highlight the novel aspects of Lux. In Section 4 we present a range of results that illustrate how Lux can precisely infer stellar labels from the APOGEE data, as well as on synthetic stellar spectra generated with Korg (Wheeler et al., 2023). In Section 5 we illustrate how Lux can effectively perform multi-survey translation between the APOGEE-GALAH surveys. We finish by discussing the idea of Lux in the wider context of stellar label determination and machine-learning in Section 6, and close by providing our summary and concluding remarks in Section 7.
@@ 3.1 Model structure
where here $\sigma_{\boldsymbol{f},n}$ represents the (APOGEE) per-pixel flux uncertainties and we include an additional variance per pixel $\boldsymbol{s}_{f}$ as a set of free parameters in the likelihood that is meant to capture the intrinsic scatter and any uncharacterized systematic errors in the spectral data (e.g., from sky lines). In principle, we could add a similar “extra variance” to the stellar labels but from experimentation we have found this to be unneeded.
@@ 3.2 Inferring parameters of a Lux model
$g(\boldsymbol{A},\boldsymbol{B},\{\boldsymbol{z}_{n}\}_{N},\boldsymbol{s}_{f})=\ln\mathcal{L}+\Omega\sum_{n}^{N}\sum_{p}^{P}z_{pn}$ (7)
@@
$g(\boldsymbol{A},\boldsymbol{B},\{\boldsymbol{z}_{n}\}_{N},\boldsymbol{s}_{f})=\sum_{n}^{N}\left[\ln p(\boldsymbol{\ell}_{n}\,|\,\boldsymbol{A},\boldsymbol{z}_{n})+\ln p(\boldsymbol{f}_{n}\,|\,\boldsymbol{B},\boldsymbol{z}_{n},\boldsymbol{s}_{f})+\Omega\sum_{p}^{P}z_{pn}\right]$ (8)
@@ 4.1 Initialization of the latent parameters
In order to optimize the parameters in Lux, $\boldsymbol{A},\boldsymbol{B},\boldsymbol{z}_{n}$, and scatter in the flux pixels, $\boldsymbol{s}$, we must first initialize them. We initialize $\boldsymbol{A}$ and $\boldsymbol{B}$ randomly from a uniform distribution over $[0,1]$ with shapes $\boldsymbol{A}=[M\times P]$ and $\boldsymbol{B}=[\Lambda\times P]$. For the latent vectors $\boldsymbol{z}_{n}$, we initialize following a similar procedure to the pre-computation of feature vectors described in the Cannon model (Ness et al., 2015). In more detail, we resize all the labels by a given centroid and scale equal to the $50^{\mathrm{th}}$ and $(97.5^{\mathrm{th}}-2.5^{\mathrm{th}})/4$ percentile values of the sample distribution, respectively [^2] We divide $(97.5^{\mathrm{th}}-2.5^{\mathrm{th}})/4$ by four as the $95^{\mathrm{th}}$ percentile range is approximately $\sim 4~{}\sigma$.. This is computed for numerical stability as some labels will have scales around $10^{3}$ whilst others may have relevant scales around $10^{-1}$; this way, all values are around unity. Using these re-scaled label values, we initialize the latent vectors for each star, $\boldsymbol{z}_{n}$, as
@@ 4.2 Training step
The training step of Lux consists of running a two-part procedure. In the first part (Agenda 1), we optimize the parameters $\boldsymbol{A},\boldsymbol{B},\boldsymbol{z}_{n}$, without any regularization. We do this using a custom multi-step optimization scheme. The first step (the $a$-step) optimizes $\boldsymbol{A}$ using the stellar label data at fixed $\boldsymbol{z}$; the second step ($b$-step) optimizes $\boldsymbol{B}$ using stellar flux data at fixed $\boldsymbol{z}_{n}$; the third step ($z$-step) then optimizes $\boldsymbol{z}_{n}$ using the stellar label and stellar spectra data at fixed (and newly optimized) $\boldsymbol{A}$ and $\boldsymbol{B}$. Here, the optimization in all three steps is performed assuming there is no scatter in the fluxes/pixels and no regularization. In the case of a linear Lux model (as we use here), these optimization steps can be done using closed-form least-squares solutions. However, for future generalizations, we instead use a Gauss–Newton least squares solver for each step (using the GaussNewton solver in JAXopt; Blondel et al. 2021). [^3] Even though it is overkill to use a nonlinear least-squares solver for this particular model form, we have found that the solutions converge very fast here because the Hessian is tractable and can be computed exactly with JAX. A run through the $a$, $b$, and $z$ steps completes one iteration. After testing different numbers of iterations and inspecting the accuracy of the model, we have found that the model reaches a plateau after five iterations [^4] Model accuracy is calculated by computing a $\chi^{2}$ metric summed over all labels, fluxes, and stars.. Thus, we run this first agenda for five iterations.
@@ 4.4 Accuracy of predicted stellar spectra
Figure 4: Reduced $\chi^{2}$ value for all stars in the validation/test set using the stellar fluxes. These stellar fluxes are generated from the latent vectors inferred from the stellar spectra themselves. $\chi^{2}_{\mathrm{r}}$ is computed as $\chi^{2}$ divided by the number of pixels in the spectrum. The majority of our test objects yield a $\chi^{2}_{\mathrm{r}}$ around unity, indicating that the Lux model is a good fit and the extent of the match between the observed (APOGEE) stellar spectra and estimates from the Lux model are in accord with the error variance.
@@ 4.8 Tests on different stellar types
As a final test of Lux model’s ability to infer stellar labels, we evaluate its performance across different stellar types. We train an Lux model using 4,000 RGB, MS, and dwarf stars from the high-SNR field all-train sample. The model predicts six labels: $T_{\mathrm{eff}}$, $\log~{}g$, [Fe/H], [Mg/Fe], $v_{\mathrm{micro}}$, and $v_{\mathrm{sin}i}$, with the latter two parameters providing additional constraints on stellar type classification. We validate the model using 1,000 stars from the high-SNR field all-test sample.
@@ 5 Results: Label transfer between APOGEE and GALAH
In detail, we identify 5,000 overlapping red giant branch stars between the APOGEE and GALAH surveys. We then divide this parent sample into a training set comprised of 4,000 stars, and a test set of 1,000 stars. As described in Section 2, we will label these as GALAH-APOGEE field giants-train and GALAH-APOGEE field giants-test, respectively.
@@
Figure 13 shows the distribution of seven stellar labels in the [X/Fe]-[Fe/H] plane for the 1,000 stars in the GALAH-APOGEE field giants-test set; here we only show the stellar labels that are not possible to be determined in APOGEE777[Na/Fe] and [Ce/Fe] has been shown to be possible to be determined in ASPCAP, albeit for a relatively small number of the APOGEE data set. Moreover, these are chemical abundance ratios that ASPCAP struggles to determine precisely due to the weak atomic lines in the $1.5-1.7~{}\mu$m regime.. We show the estimated Lux labels in the top row, and the corresponding GALAH labels in the bottom row. Here, the Lux labels are determined by optimizing the latent representations for the test set stars using each star’s spectral fluxes, and projecting those to stellar labels using $\boldsymbol{A}$ through Equation 1. We find that Lux model is able to infer stellar labels determined in GALAH for APOGEE stars, and in some cases, is able to estimate an abundance for stars that do not report a GALAH stellar label [^8] These are the stars that appear as a horizontal stripe in Figure 13, which have been set to the median value of the distribution by our model due to GALAH not being able to provide a measurement.; this is one of the main advantages of Lux framework, as it is able to operate with partial missing labels. The full test set validation is shown in Figure 17 in Appendix B.
@@ 6.2.1 Stellar label inference
Lux can be used to infer stellar parameters from spectroscopic survey data by learning from a training set with known parameters. This is valuable for efficiently determining parameters for large stellar surveys (e.g., SDSS, Gaia, LAMOST, GALAH, DESI, 4MOST, WEAVE) by emulating more costly pipeline runs. One immediate application is determining stellar parameters and abundances for stars in the SDSS-V Galactic Genesis Survey (Kollmeier et al., 2017), which is collecting millions of APOGEE spectra. Lux could also be used to determine spectro-photometric distances from a reliable training set (Hogg et al., 2019), or compile catalogs of stellar ages for giants stars based on [C/N] abundances and asteroseismology (Ness & Lang, 2016).
@@ 6.2.2 Multi-survey translation
Lux enables translation between the notoriously different stellar parameter outputs of different surveys and instruments by training on overlapping sources. This allows determination of parameters that may be difficult or impossible to measure directly in one survey but are well-measured in another. For example, stellar parameters and abundances could potentially be determined for the vast set of Gaia XP spectra by training on stars that overlap with APOGEE (e.g., Andrae et al., 2023a; Li et al., 2023). Similarly, APOGEE-quality stellar labels could potentially be determined for BOSS spectra using overlapping stars from SDSS-V Milky Way Mapper. However, care must be taken to validate that the translated parameters reflect genuine spectral features rather than just correlations in the training set.
@@ Acknowledgements
The authors would like to thank Adam Wheeler for providing the Korg spectra, Julianne Dalcanton for enlightening conversations about future prospects of Lux, Carrie Filion for all the help and support, Catherine Manea, David Nidever, Andrew Saydjari, Greg Green, Hans-Walter Rix, and the CCA stellar spectroscopy, CCA Astronomical Data, and CCA Nearby Universe groups for helpful discussions. DH would also like to thank Sue, Alex, and Debra for everything they do. The Flatiron Institute is a division of the Simons Foundation.

### v2

@@ Abstract
The large volume of spectroscopic data available now and from near-future surveys will enable high-dimensional measurements of stellar parameters and properties. Current methods for determining stellar labels from spectra use physics-driven models, which are computationally expensive and have limitations in their accuracy due to simplifications. While machine learning methods provide efficient paths toward emulating physics-based pipelines, they often do not properly account for uncertainties and have complex model structure, both of which can lead to biases and inaccurate label inference. Here we present Lux: a data-driven framework for modeling stellar spectra and labels that addresses prior limitations. Lux is a generative, multi-output, latent variable model framework built on JAX for computational efficiency and flexibility. As a generative model, Lux properly accounts for uncertainties and missing data in the input stellar labels and spectral data and can either be used in probabilistic or discriminative settings. Here, we present several examples of how Lux can successfully emulate methods for precise stellar label determinations for stars ranging in stellar type and signal-to-noise from the APOGEE survey. We also show how a simple Lux model is successful at performing label transfer between the APOGEE and GALAH surveys. Lux is a powerful new framework for the analysis of large-scale spectroscopic survey data. Its ability to handle uncertainties while maintaining high precision makes it particularly valuable for stellar survey label inference and cross-survey analysis, and the flexible model structure allows for easy extension to other data types.
@@ 1 Introduction
In an attempt to circumvent these requirements, in the last decade there has been a push to use data-driven methods to determine stellar parameters and element abundances of stars using linear regression (e.g., Ness et al., 2015; Casey et al., 2016) or machine-learning methods (e.g., Ting et al., 2019; Ciuca & Ting, 2022; Andrae et al., 2023a; Buck & Schwarz, 2024; Różański et al., 2024), that are more suited to deal with high-volume data. These methods are considered latent variable models, that aim to infer a mapping between observables and measurables via some latent space (e.g., Nolan et al., 2006; Hall et al., 2019; Eilers et al., 2019), and can fall under the umbrella of “emulator” or “label transfer” approaches depending on the training and testing data employed. In essence, they are used to: 1) train a model on a set of (trustworthy) input stellar spectra and stellar labels; 2) optimize a set of (latent) model parameters; 3) use the trained model to predict stellar labels for some catalog data. While the functional form of the model may vary between these approaches (e.g., quadratic, neural-network, etc…), the process is still the same in practice. These models have proven extremely successful in delivering accurate and precise stellar labels in a fast and cost-effective manner (e.g., Ness et al., 2016; Ho et al., 2017b, a; Xiang et al., 2017, 2019; Andrae et al., 2023b; Li et al., 2024; Guiglion et al., 2024; Ness et al., 2024). However, they also come with some limitations. For example, these models assume the input training stellar labels are ground-truth (i.e., no uncertainties are taken into account), and are not able to train a model with missing label data (for example, stars that have $T_{\mathrm{eff}}$ information but no [Fe/H]), although see Leung & Bovy (2018) or Koblischke & Bovy (2024) for counter examples. As a result, many good data are not used in the training step, which hinders the stellar label regime that can be probed. This restriction also limits the ability to perform a two-way label transfer between two spectroscopic data sets, as typically stars will have a set of labels from one survey but not the other.
@@
In Section 2 we introduce the data and samples used. In Section 3 we describe the framework of the Lux model, introduce the likelihood function, explain the routine for employing the model, and highlight the novel aspects of Lux. In Section 4 we present a range of results that illustrate how Lux can precisely infer stellar labels from the APOGEE data. We also show results of the Lux mode using synthetic stellar spectra generated with Korg (Wheeler et al., 2023) in Appendix C. In Section 5 we illustrate how Lux can effectively perform multi-survey translation between the APOGEE-GALAH surveys. We finish by discussing the idea of Lux in the wider context of stellar label determination and machine-learning in Section 6, and close by providing our summary and concluding remarks in Section 7.
@@ 3.1 Model structure
where here $\sigma_{\boldsymbol{f},n}$ represents the (APOGEE) per-pixel flux uncertainties and we include an additional variance per pixel $\boldsymbol{s}_{f}$ as a set of free parameters in the likelihood that is meant to capture the intrinsic scatter and any uncharacterized systematic errors in the spectral data (e.g., from sky lines). In principle, we could add a similar “extra variance” to the stellar labels but from experimentation we have found this to be unneeded. However, the Lux software has been developed to be able to include per label scatter terms if wanted.
@@ 3.2 Inferring parameters of a Lux model
$g(\boldsymbol{A},\boldsymbol{B},\{\boldsymbol{z}_{n}\}_{N},\boldsymbol{s}_{f})=\ln\mathcal{L}+\Omega\sum_{n}^{N}\sum_{p}^{P}z_{pn}^{2}$ (7)
@@
$g(\boldsymbol{A},\boldsymbol{B},\{\boldsymbol{z}_{n}\}_{N},\boldsymbol{s}_{f})=\sum_{n}^{N}\left[\ln p(\boldsymbol{\ell}_{n}\,|\,\boldsymbol{A},\boldsymbol{z}_{n})+\ln p(\boldsymbol{f}_{n}\,|\,\boldsymbol{B},\boldsymbol{z}_{n},\boldsymbol{s}_{f})+\Omega\sum_{p}^{P}z_{pn}^{2}\right]$ (8)
@@ 4.1 Initialization of the latent parameters
In order to optimize the parameters in Lux, $\boldsymbol{A},\boldsymbol{B},\boldsymbol{z}_{n}$, and scatter in the flux pixels, $\boldsymbol{s}$, we must first initialize them. We initialize $\boldsymbol{A}$ and $\boldsymbol{B}$ randomly from a uniform distribution over $[0,1]$ with shapes $\boldsymbol{A}=[M\times P]$ and $\boldsymbol{B}=[\Lambda\times P]$. For the latent vectors $\boldsymbol{z}_{n}$, we initialize following a similar procedure to the pre-computation of feature vectors described in the Cannon model (Ness et al., 2015). In more detail, we resize all the labels by a given centroid and scale equal to the $50^{\mathrm{th}}$ and $(97.5^{\mathrm{th}}-2.5^{\mathrm{th}})/4$ percentile values of the sample distribution, respectively [^2] We divide $(97.5^{\mathrm{th}}-2.5^{\mathrm{th}})/4$ by four as the $95^{\mathrm{th}}$ percentile range is approximately $\sim 4~{}\sigma$.. This is appropriate for two main reasons: $i$) it is safe numerically, as some labels may have scales around $10^{3}$ whilst others may have relevant scales around $10^{-1}$. This way, all values are around unity; $ii$) because we are assuming a linear relationship between (functions of) the stellar labels and the spectral fluxes (as done in the Cannon (Ness et al., 2015) that is very successful in these kinds of tasks). Using these re-scaled label values, we initialize the latent vectors for each star, $\boldsymbol{z}_{n}$, as
@@ 4.2 Training step
The training step of Lux consists of running a two-part procedure. In the first part (Agenda 1), we optimize the parameters $\boldsymbol{A},\boldsymbol{B},\boldsymbol{z}_{n}$, without any regularization. We do this using a custom multi-step optimization scheme. The first step (the $a$-step) optimizes $\boldsymbol{A}$ using the stellar label data at fixed $\boldsymbol{z}_{n}$; the second step ($b$-step) optimizes $\boldsymbol{B}$ using stellar flux data at fixed $\boldsymbol{z}_{n}$; the third step ($z$-step) then optimizes $\boldsymbol{z}_{n}$ using the stellar label and stellar spectra data at fixed (and newly optimized) $\boldsymbol{A}$ and $\boldsymbol{B}$. Here, the optimization in all three steps is performed assuming there is no scatter in the fluxes/pixels and no regularization. In the case of a linear Lux model (as we use here), these optimization steps can be done using closed-form least-squares solutions. However, for future generalizations, we instead use a Gauss–Newton least squares solver for each step (using the GaussNewton solver in JAXopt; Blondel et al. 2021). [^3] Even though it is overkill to use a nonlinear least-squares solver for this particular model form, we have found that the solutions converge very fast here because the Hessian is tractable and can be computed exactly with JAX. A run through the $a$, $b$, and $z$ steps completes one iteration. After testing different numbers of iterations and inspecting the accuracy of the model, we have found that the model reaches a plateau after five iterations [^4] Model accuracy is calculated by computing a $\chi^{2}$ metric summed over all labels, fluxes, and stars.. Thus, we run this first agenda for five iterations.
@@ 4.4 Accuracy of predicted stellar spectra
Figure 4: Reduced $\chi^{2}$ value for all stars in the validation/test set using the stellar fluxes. The median value (dotted line) is 1.07. These stellar fluxes are generated from the latent vectors inferred from the stellar spectra themselves. $\chi^{2}_{\mathrm{r}}$ is computed as $\chi^{2}$ divided by the number of pixels in the spectrum. The majority of our test objects yield a $\chi^{2}_{\mathrm{r}}$ around unity, indicating that the Lux model is a good fit and the extent of the match between the observed (APOGEE) stellar spectra and estimates from the Lux model are in accord with the error variance.
@@ 4.8 Tests on different stellar types
As a final test of Lux model’s ability to infer stellar labels, we evaluate its performance across different stellar types. We train an Lux model using 4,000 RGB, MS, and dwarf stars from the high-SNR field all-train sample. The model predicts six labels: $T_{\mathrm{eff}}$, $\log~{}g$, [Fe/H], [Mg/Fe], $v_{\mathrm{micro}}$, and $v_{\mathrm{sin}i}$, with the latter two parameters providing additional constraints on stellar type classification [^7] We note that it is not vital to include these last two parameters for distinguishing stellar types, as $\log~{}g$ should suffice to distinguish e.g., giant stars from dwarfs. However, during testing we found that including $v_{\mathrm{micro}}$ and $v_{\mathrm{sin}i}$ did help increase label precision.. We validate the model using 1,000 stars from the high-SNR field all-test sample.
@@ 5 Results: Label transfer between APOGEE and GALAH
In detail, we identify 5,000 overlapping red giant branch stars between the APOGEE and GALAH surveys. We then divide this parent sample into a training set comprised of 4,000 stars, and a validation set of 1,000 stars. As described in Section 2, we will label these as GALAH-APOGEE field giants-train and GALAH-APOGEE field giants-test, respectively.
@@
Figure 13 shows the distribution of seven stellar labels in the [X/Fe]-[Fe/H] plane for the 1,000 stars in the GALAH-APOGEE field giants-test set; here we only show the stellar labels that are not possible to be determined in APOGEE888[Na/Fe] and [Ce/Fe] has been shown to be possible to be determined in ASPCAP, albeit for a relatively small number of the APOGEE data set. Moreover, these are chemical abundance ratios that ASPCAP struggles to determine precisely due to the weak atomic lines in the $1.5-1.7~{}\mu$m regime.. We show the estimated Lux labels in the top row, and the corresponding GALAH labels in the bottom row. Here, the Lux labels are determined by optimizing the latent representations for the test set stars using each star’s spectral fluxes, and projecting those to stellar labels using $\boldsymbol{A}$ through Equation 1. We find that Lux model is able to infer stellar labels determined in GALAH for APOGEE stars, and in some cases, is able to estimate an abundance for stars that do not report a GALAH stellar label [^9] These are the stars that appear as a horizontal stripe in Figure 13, which have been set to the median value of the distribution by our model due to GALAH not being able to provide a measurement.; this is one of the main advantages of Lux framework, as it is able to operate with partial missing labels. The full test set validation is shown in Figure 17 in Appendix B.
@@ 6.2.1 Stellar label inference
Lux can be used to infer stellar parameters from spectroscopic survey data by learning from a training set with known parameters. This is valuable for efficiently determining parameters for large stellar surveys (e.g., SDSS: Blanton et al., 2017; Kollmeier et al., 2017, Gaia: Gaia Collaboration et al., 2020, LAMOST: Zhao et al., 2012, GALAH: Freeman, 2012, DESI: Cooper et al., 2023, 4MOST: de Jong et al., 2012, WEAVE: Jin et al., 2023) by emulating more costly pipeline runs. One immediate application is determining stellar parameters and abundances for stars in the SDSS-V Galactic Genesis Survey (Kollmeier et al., 2017), which is collecting millions of APOGEE spectra. Lux could also be used to determine spectro-photometric distances from a reliable training set (Hogg et al., 2019), or compile catalogs of stellar ages for giants stars based on [C/N] abundances and asteroseismology (Ness & Lang, 2016).
@@ 6.2.2 Multi-survey translation
Lux enables translation between the notoriously different stellar parameter outputs of different surveys and instruments by training on overlapping sources. This allows determination of parameters that may be difficult or impossible to measure directly in one survey but are well-measured in another. For example, stellar parameters and abundances could potentially be determined for the vast set of Gaia XP spectra by training on stars that overlap with APOGEE (e.g., Andrae et al., 2023a; Li et al., 2024). Similarly, APOGEE-quality stellar labels could potentially be determined for BOSS spectra using overlapping stars from SDSS-V Milky Way Mapper. However, care must be taken to validate that the translated parameters reflect genuine spectral features rather than just correlations in the training set.
@@ Acknowledgements
The authors would like to thank Adam Wheeler for providing the Korg spectra, Julianne Dalcanton for enlightening conversations about future prospects of Lux, Carrie Filion for all the help and support, Catherine Manea, David Nidever, Andrew Saydjari, Greg Green, Hans-Walter Rix, and the CCA stellar spectroscopy, CCA Astronomical Data, and CCA Nearby Universe groups for helpful discussions. The authors would also like to thank the anonymous reviewer for a helpful report on earlier versions of this manuscript. DH would also like to thank Sue, Alex, and Debra for everything they do. The Flatiron Institute is a division of the Simons Foundation.